OpenStudy (anonymous):

What are the roots of the quadratic equation (x-1)(4-x)=2

OpenStudy (tkhunny):

That '2' is no good. There must be '0' over there. Make it so.

OpenStudy (anonymous):

on simplifying 4x-x^2-4+x=2 -----> x^2-5x+6=0 ---->(x-3).(x-2)=0 x=2 and x=3

OpenStudy (radar):

Do you need to see all the steps of the simplication process?

OpenStudy (anonymous):

Yeah

OpenStudy (mathmale):

Hey, Craig, what HunnyHunk is driving at, in more formal language, is that we need to re-write the given quadratic equation in standard form before attempting to solve it. Start with (x-1)(4-x)=2. Multiply the two binomials together. Then re-arrange all the terms on the left side according to decreasing powers of x.

OpenStudy (mathmale):

A quadratic equation in standard form looks like ax^2 + bx + c = 0.

OpenStudy (mathmale):

Note that @Rohitkhanna has already done exactly what I am recommending.

OpenStudy (radar):